20130905 Instrument Initial Night-time Checkout

Configured for MODS at start of night

During the afternoon, John M and Jenny swapped the sequencer card and fiber board from M1.RC and installed these in the spare IC in CRB. Jerry rigged this spare IC as M1.RC. The original M1.RC will be shipped to OSU as soon as possible. This swap appears to have resolved the readout problem seen earlier in 3k x 3k and 4k x 3k mode.

Before opening ran started up MODS: mods1x.20130905.00NN.fits:

b r  
1,2 1,2 sieveSnap
3 3 imSnap
4-10 4-10 grlamps and prlamps, but telescope was not stationary during these. Will repeat.

02:45 UT initial preset. No star found. Required an offsets of -383" IE, -37" CA to find an ACT star! Investigating this:
  • M2 had been homed. But David is homing M2 from OSS again. Once homed, star is back to center, so that was not the problem.
  • David stopping/restarting PCS. M2 initialized on PSF also. Pointing model offsets reset. No star seen, still. Encoders were reset, but...
  • 03:20 David will restart mcspu. Still no star nearby.
  • 03:30 David reset encoders. John H finds that gps is not locked (on MCS Power Drive Status gui). Previous star for which deltaIE = -383 was at elev=86, az=80. Try a star near meridian and dec=25. Cannot find star.
  • 03:30-04:00. Checking GPS in dome. Kevin, Dan Cox on phone. The GPSLOCK off on MCS Power Drive Status may be a 'red herring'.
  • 04:00 Send active preset to Polaris, using a MODS script (in /home/lbto/SciPrep/MODS called Polaris.img). Guide probe on-axis. Star not seen and shell ripped.
Thu Sep 5 04:06:54.692 2013 Preset alert for left MODS mode ACTIVE: AOSL
Thu Sep 5 04:06:54.715 2013 Preset alert for left MODS mode ACTIVE: Stop[397] complete
  • From MODS gui, I "homed" the guide probe. Will resend the preset once shell is recovered. Some problems recovering shell and Doug M was called. Doug resetting shell. In meantime, sending an "acquire" preset to Polaris. ptreset. Polaris found. Time may be a problem. Going back to a pointing star...near Polaris it is OK. Try a pointing star at lower declination. We see the star. It seems that homing the MODS guide probe fixed it. Perhaps the 383" arcsec is from another M5 star??? But so close together? Shell is fixed now.

04:38 David collimating.

04:40 Measure center of rotation on guider.
  • point to BS 9177 with gs=0 (guider on-axis). Collimate.
  • David moved rotator to 0 deg.
  • I moved guide probe on-axis "modsCmd gprobe 0 0"
  • cd /home/lbto/20130905/ and take an exposure "readGuideCam left -e 30000"
    • first exposure after guiding typically times out (only happens with guider, not WFS). AzCam server timed out waiting for full CCD readout, guide image invalid.
    • first real exposure is left_guider0000001.fits. Star at (234, 178).
    • take image while rotator slewing:
      • David prep command "rslewtohold ldg 360"
      • I issue "readGuideCam left -e 80000".
      • 2 seconds after image started, David issues "rslewtohold".
      • result is 000002.
        • center of rotation (269, 199)
    • recollimate. Stop guiding thread.
      • 000003 is with star in beamsplitter
      • 000004 is after issuing modsCmd gprobe 0 0 (229, 198)
      • 000005 with rotator slewing. (271, 203)

John's analysis of the MODS1 rotator axis on the guider

What happens now (since MODS1 commissioning) using MODS_L.cfg:

1) Between the 1st and 2nd acquisition images, the telescope is offset
to position the star at unbinned untrimmed CCD pixel 258,199
(guidecam_center_x & y).  The MODS1 acquisition images have 24 pixels
trimmed from X, so CCD 258,199 appears on the image (ds9 coords) at 234,199.

2) Between the 2nd and 3rd acquisition images, the agw stage (not the
telescope) is offset to move the star from CCD 258,199 pixels
(guidecam_center_x & y) to CCD 249,405 pixels (guidecam_hotspot_x & y).
This probe offset is +0.37,-13.1 mm (WFScam_offset_x & y)
In image coordinates, the guidecam_hotspot appears at 225,405 pixels.

The problem:
The small issue is that the rotator center appears on the MODS1 guider
at 271,203 +-2 pixels in image coordinates and CCD 295,203 pixels.
The fact that we are not guiding at the rotator center on axis
introduces a small epicycle in the position of the wfs pupil on the
lenslets as the rotator angle changes.  The simple relation of X-Y
position of the guide star in the field to pupil position now has a
rotator angle term.

We (Doug M and John H) are proposing to adjust the MODS1 on-axis
guiding position to the rotator center to remove this source of noise
in the pupil position and the wfs results.

The plan:
Adjust guidecam_center_x & y so that the acquired star is placed on
the rotator axis at CCD 295,203 pixels (increasing the X-value by 37
pixels.

Simultaneously counter-adjust the guide probe offset so the star still
lands on the guiding pinhole when the probe moves.  That means
increasing X-offset by:
37 pixels * 0.106 arcsec/pixel * 0.600 mm/arcsec = 2.35 mm
adding the old offset 0.37 + 2.35 = 2.72 mm

(This calculation ignores any rotational skew of the axes - which
looks to be about 0.2 mm / 13.1.)

The new probe offset should be 2.72,13.1 mm.

05:18 Slit scan
  • Executing script /home/lbto/SciPrep/MODS/Slitscan/StoneOgs1344_slit.img
  • Seeing is 1.2-1.5". We are at airmass 1.213.
  • Collimated with ACTIVE preset. Changed script to send TRACK PRESET (and remove WFSWait) and make slitscan. Current LEFTZEROPOINT is 27.63.
  • red images 12-16, 17 is slit image.
  • note that the stars have tails, even on-axis.
#   COL    LINE     COORDINATES R    MAG           FLUX     SKY    PEAK    E     PA BETA ENCLOSED   MOFFAT DIRECT
1722.50 1797.64 1722.50 1797.64 30.45  11.04 383777.   2378.   2250. 0.19   88 2.14    10.48    10.02  10.17
1722.87 1674.18 1722.87 1674.18 29.22  11.01 392792.   2386.   2628. 0.20   86 2.59     9.85     9.51   9.75
1718.02 1551.90 1718.02 1551.90 34.29  11.91 171509.   2364.   2314. INDEF    0 4.51    11.01    10.56  13.01
1720.14 1428.52 1720.14 1428.52 23.54  10.93 425367.   2134.   3670. 0.18   88 2.87     8.20     8.12   7.84
1720.91 1305.63 1720.91 1305.63 26.61  10.87 447615.   2120.   3550. 0.18  -89 3.05     8.33     8.44   9.00
  • line drawn between top and bottom star image is slightly tilted with respect to slit, but it may be influenced by drift.

05:42 Repeating slit scan, this time on StoneO but with a guide star that is in the MODS field
  • StoneOgs1224_slit_active.img Seeing~1.1-1.2". Took a few cycles to collimate.
  • StoneOgs1224_slit_track.img
  • 18-23, 24 is slit

06:04 Repeating slit scan, now bouncing between top/bottom of slit central region a few times. StoneOgs1224_slit_bounce.img
  • Seeing~1.4-1.6".
  • Note that stars look much rounder. I did not look at 18-23 but suspect these too are rounder than the first set, since the guide star was in the MODS field and not out.

06:33 Check AGw transform by sending an active preset to the same field as above, and obtaining a set of images with relative offsets between.
  • cd /home/lbto/SciPrep/MODS/AlignCheck
  • execMODS StoneOgs1224_dice5.img
  • seeing is bouncing up to 2.5"

07:05 John edited the MODS_L.cfg (see his "plan" above) to change guidecam_center_x from 258 to 295, WFScam_offset_x from 0.37 to 2.72. David stopped/restarted GCS. ACTIVE on-axis preset to BS 9188. These changes were correct and cause the guide star to be centered on the rotator center.

07:20 Checked seeing closer to zenith, very poor still. Stay with MODS.

07:25 Check Rotator Center.
  • /home/lbto/SciPrep/RotatorCenter/
    • execMODS StoneP1_act0.img to collimate b=15, r=42
    • execMODS StoneP1_trk0.img then StoneP1_trk180.img to obtain 30-sec images of the field with PA=0 and PA=180.
    • trk0: b=17, r=44
    • trk180: b=18,r=45
    • back to collimate
    • trk90: b=20, r=47
    • trk270: b=21,r=48

08:00 acqMODS ~modseng/modsScripts/modsSpecPhot/feige110.acq
  • =nb: using CISCO VPN to obs3, need to open a ds9 window before running modsAlign, otherwise modsAlign can fail with a race condition.
    • modsAlign -r mods1r.20130905.0049.fits
    • centered pretty well (0050). Repeated the thru-slit since seeing is poor & variable and guide star is bouncing around (0051). Seeing 1.5-2.3".
    • execMODS ~modseng/modsScripts/modsSpecPhot/feige110.obs

**Michelle starts driving**

08:32 * ERROR: Command 'instconfig blue grating' timed out after 60 seconds Abort, Retry, or Ignore? * ERROR: instconfig instrument configuration already in progress - command disallowed * Refresh on GUI fixed the problem

09:00 acqMODS ~modseng/modsScripts/modsSpecPhot/g191b2b.acq
  • modsAlign -r mods1r.20130905.0064.fits
  • not centered well (0065). Applied a further -1.5 offset in X.(0066). Ok a little overzealous with the offset; +0.5 in X (0067). Seeing 1.3-1.7".
  • execMODS ~modseng/modsScripts/modsSpecPhot/g191b2b.obs

09:30 acqMODS ~modseng/modsScripts/modsSpecPhot/wolf1346.acq
  • modsAlign -r mods1r.20130905.0071.fits
  • not too bad (0072). Applied offset -0.5 in X. Seeing 1.5-1.7" -- but high airmass and lots of coma (?) in GCS images
  • execMODS ~modseng/modsScripts/modsSpecPhot/wolf1346.obs

09:55 acqMODS ~modseng/modsScripts/modsSpecPhot/gd71.acq
  • =preset failed!
  • tried again; then tried a few pointings at bright stars -- no luck
  • looks like lots of clouds are moving in!

10:05 execMODS MODSsleep.pro

10:10 Switching to LBC

10:23 Running test bias on LBC and setting up

10:33 First offset to see if we see any stars; Yes!

10:40 Running dofpia on WT10_242. Mirror temp SX = 11.0; DX = 10.7. Ambient 10.2. DIMM seeing = 1.2-1.5"

  • Converged but only with 3-4 pupils. Central holes are visable though. Headed to another field to see if I can get more pupils

10:50 Re-run dofpia on a new field (WT10_261). Many more pupils. Again, at first blush, convergence is good. Seeing value reported is good. Also completed co-pointing on target.

Btimesta= 105524 Rtimesta= 105521

Old Delta New ------- ------- ------- CE -37.608 8.257 -29.351 CA -109.506 11.626 -97.880

11:00 Slew to target NGC110 @ El = 46; Az = 340 @ 1.38 airmass

* Dithering and guiding basic checks. Overall stars look; maybe some elongation in blue but seeing isn't good enough to tell on most images. So far, guiding gives no obvious errors and seems to run for both 60s and 180s exposures and dithering is working as expected, although I'll carefully check

11:36 dofpia

11:37 seeing has improved enough (~1.0-1.2"") that I think I can do a quick check of the rotation center

  • running rotatorcenter_pa0 and rotatorcenter_pa180 on NGC110.ob
    • lbcr.20130905.114105.fits = rotatorcenter_pa0
    • lbcb.20130905.114106.fits = rotatorcenter_pa0
    • lbcr.20130905.114437.fits = rotatorcenter_pa180
    • lbcb.20130905.114440.fits = rotatorcenter pa_180

11:42 dofpia

11:50 Second pair of rotatorcenter, as suggested:

  • lbcr.20130905.115216.fitss = rotatorcenter_pa0
  • lbcb.20130905.115220.fits = rotatorcenter_pa0
  • lbcr.20130905.115418.fits = rotatorcenter_pa180
  • lbcb.20130905.115421.fits = rotatorcenter pa_180

ROTCEN
   RED   1081.0, 2913.7  and  1079.7,2914.3  (old: 1078. 2913.)    USE: 1080,2914
   BLUE  1031.7, 2921.9  and  1031.2,2921.7  (old: 1035. 2924.)    USE: 1031,2922

12:00 Seeing has degraded -- around 2" on the images. DIMM is reading 1.3; may be target is too low. Trying one last target ACT0342 @ X = 1.00

12:05 dofpia

  • DIMM is reporting 1.7" at X = 1.00. Too bad for anymore LBC tests tonight. Closing!

-- OlgaKuhn - 05 Sep 2013
Topic revision: r6 - 06 Sep 2013, DavidThompson
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback